# Driving Beyond Privilege: Distilling Dense-Reward Knowledge into Sparse-Reward Policies

### v1

@@ Abstract
We study how to exploit dense, simulator-defined rewards in vision-based autonomous driving without inheriting their misalignment with deployment metrics. In realistic simulators such as CARLA, privileged state (e.g., lane geometry, infractions, time-to-collision) can be converted into dense rewards that greatly stabilize and accelerate model-based reinforcement learning, but policies trained directly on these shaped signals often overfit and fail to generalize when evaluated on sparse task objectives such as route completion and collision-free overtaking. We propose reward-privileged world-model distillation, a two-stage framework in which a teacher DreamerV3-style agent is first trained with a dense, privileged reward, and only its latent dynamics are subsequently distilled into a student trained solely on sparse task rewards. Teacher and student share the same observation space (semantic bird’s-eye-view images); privileged information enters only through the teacher’s reward, and the student does not imitate the teacher’s actions or value estimates. Instead, the student’s world model is regularized to match the teacher’s latent dynamics while its policy is learned from scratch on sparse success/failure signals. In CARLA lane-following and overtaking benchmarks, our sparse-reward students consistently outperform both dense-reward teachers and sparse-from-scratch baselines. On unseen lane-following routes, reward-privileged distillation improves success by approximately 23% relative to the dense teacher while maintaining comparable or better safety metrics. On overtaking, students retain near-perfect performance on training routes and achieve up to a 27× relative improvement in success on unseen routes compared to the dense teacher, with improved lane keeping. These results indicate that dense rewards can be safely leveraged to learn richer dynamics models while keeping the deployed policy optimized strictly for sparse, deployment-aligned objectives.
@@ II Related Work
Another line blends offline data and world models. World on Rails learns a global world model from pre-recorded driving logs (with access to privileged log information such as maps and other vehicles’ trajectories), performs dynamic programming in this model to compute action-values, and uses these values to supervise a reactive image-only policy [chen_learning_2021]. The resulting policy, which never sees privileged inputs at execution time, outperforms both imitation learning and standard RL on CARLA benchmarks. Overall, these works show that world models, often aided by privileged inputs, are powerful tools for autonomous driving. However, the teacher and student (or planner and policy) typically share the same underlying reward or driving score. Complementary evidence from off-road settings indicates that seemingly low-level perception design choices such as camera region-of-interest and dataset size can strongly affect the performance and robustness of end-to-end controllers [khanzada_analytical_2024], underscoring the importance of representation choices alongside the RL algorithm.

### v2

@@ Abstract
We study how to exploit dense simulator rewards in vision-based autonomous driving without inheriting their misalignment with deployment metrics. In realistic simulators such as CARLA, privileged state (e.g., lane geometry, infractions, time-to-collision) can be converted into dense rewards that greatly stabilize and accelerate model-based reinforcement learning, but policies trained directly on these signals often overfit and fail to generalize when evaluated on sparse objectives such as route completion and collision-free overtaking. We propose reward-privileged world model distillation, a two-stage framework in which a teacher DreamerV3-style agent is first trained with a dense privileged reward, and only its latent dynamics are subsequently distilled into a student trained solely on sparse task rewards. Teacher and student share the same observation space (semantic bird’s-eye-view images); privileged information enters only through the teacher’s reward, and the student does not imitate the teacher’s actions or value estimates. Instead, the student’s world model is regularized to match the teacher’s latent dynamics while its policy is learned from scratch on sparse success/failure signals. In CARLA lane-following and overtaking benchmarks, sparse-reward students consistently outperform both dense-reward teachers and sparse-from-scratch baselines. On unseen lane-following routes, reward-privileged distillation improves success by approximately 23% relative to the dense teacher while maintaining comparable or better safety. On overtaking, students retain near-perfect performance on training routes and achieve up to a 27× relative improvement in success on unseen routes, with improved lane keeping. These results show that dense rewards can be safely leveraged to learn richer dynamics models while keeping the deployed policy optimized strictly for sparse, deployment-aligned objectives.
@@ II Related Work
Another line blends offline data and world models. World on Rails learns a global world model from pre-recorded driving logs (with access to privileged log information such as maps and other vehicles’ trajectories), performs dynamic programming in this model to compute action-values, and uses these values to supervise a reactive image-only policy [chen_learning_2021]. The resulting policy, which never sees privileged inputs at execution time, outperforms both imitation learning and standard RL on CARLA benchmarks. Overall, these works show that world models, often aided by privileged inputs, are powerful tools for autonomous driving. However, the teacher and student (or planner and policy) typically share the same underlying reward or driving score. Complementary evidence from off-road settings indicates that seemingly low-level perception design choices such as camera region-of-interest and dataset size can strongly affect the performance and robustness of end-to-end controllers [khanzada2024analytical], underscoring the importance of representation choices alongside the RL algorithm.